a cylindrical metal has a height of 0.27 m and a radius of 0.11 m the electrical field is directed upward along the entire surface of the cake and including the top and bottom which with uniform magnitude of 4.0 x 10-5 N/C how many charge does the can contain


1
Expert's answer
2021-04-05T10:37:55-0400
"V=3.14\times 0.11^2\times0.27=0.0103 m^3"

"\rho =\frac {4\times 10^{-5}}{1.03\times 10^{-2}}=3.88\times 10^{-2}"